Calculate the volume (in mL) of a 0.235 M HBr solution that would react with 0.362 g KOH according to the equation,

HBr + KOH → KBr + H2O
1
Expert's answer
2017-07-28T15:44:07-0400
1) m(KOH) = 0,362g CM = n/V; V = n/CM n(KOH) = m/M = 0,362g /56g/mol =
CM(HBr) = 0,235 M =0,00646 mol; n(HBr) =0,00646 mol ;
V(HBr) - ? V(HBr) = 0,00646 mol/0,235mol/L = 0,0275L = 27,5mL
Answer: V(HBr) = 27,5mL